if your father has a savings of Rs 2,000,000/- you have a choice between spending the money now or putting it in a bank account for 5 years that pays you 5% per annum compound interest.Calculate the opportunity cost of spending money now?


1
Expert's answer
2021-04-07T10:26:04-0400

"A=P(1+\frac{r}{n})^{nt}"


Where

A=final amount

P=initial principal balance

r=interest rate

n=number of times interest applied per time period

t=number of time periods elapsed



First, convert R as a percent to r as a decimal

r = R/100

r = 5/100

r = 0.05 rate per year,



"A=2,000,000(1+\frac{0.05}{1})^{(1)(5)}=2,552,563.125"


Opportunity cost = "2,552,563.125-2,000,000=552,563.125"

Opportunity cost = Rs 552,563.125